OpenStudy (confusionist):

The volume of a cube, in cm^3 is numerically equal to its surface area in cm^2. Calculate the length of a side of the cube.

OpenStudy (anonymous):

6x^2 = x^3. Solve for positive x

OpenStudy (confusionist):

Yeah, that;s where I got. I'm having trouble figuring it out from there.

OpenStudy (confusionist):

Thanks! :) This was really helpful!

OpenStudy (anonymous):

0 = x^3 - 6x^2 = x (x^2 - 6) = x (x - sqrt(6)) (x + sqrt(6)) x = sqrt(6)